40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 21 00:45:30 2015, Koji, Update, IOO, IMC fix PDRFsignal_cable.JPGQsignal_cable1.JPGQsignal_cable2.JPGQsignal_cable3.JPG
    Reply  Sat Nov 21 00:46:33 2015, Koji, Update, IOO, IMC OLTF IMC_OLTF.pdfPMC_noise_comparison.pdf
       Reply  Sun Nov 22 12:12:17 2015, Koji, Update, IOO, IMC OLTF 
    Reply  Sun Nov 22 07:09:01 2015, rana, Update, IOO, IMC fix 
       Reply  Sun Nov 22 12:07:09 2015, Koji, Update, IOO, IMC fix 
          Reply  Mon Nov 23 22:12:10 2015, Koji, Update, IOO, LO level check for the IMC demod board IMG_2137.JPGIMG_2138.JPG
             Reply  Wed Nov 25 04:24:21 2015, rana, Update, IOO, LO level check for the IMC demod board 
                Reply  Wed Nov 25 10:07:10 2015, Koji, Update, IOO, LO level check for the IMC demod board 
                   Reply  Wed Nov 25 16:46:10 2015, Koji, Update, IOO, LO level check for the IMC demod board D990511-40m_151123.pdf
                      Reply  Wed Nov 25 20:07:35 2015, rana, Update, IOO, LO level check for the IMC demod board 
                   Reply  Mon Nov 30 14:12:14 2015, ericq, Update, LSC, LO level check for the LSC RF distribution box 
                      Reply  Mon Nov 30 15:17:57 2015, Koji, Update, LSC, LO level check for the LSC RF distribution box 
Message ID: 11794     Entry time: Sat Nov 21 00:45:30 2015     Reply to this: 11795   11796
Author: Koji 
Type: Update 
Category: IOO 
Subject: IMC fix 

Based on the observation of the PMC error signal, I started measuring the IMC OLTF. Immediately, it was found that the overall IMC loop gain was too low.
The UGF was ~40kHz, which was really marginal. It had been >100kHz when I have adjusted it about a year ago. (Next entry for the detail)

The first obvious thing was that the SMA cables around the IMC servo have visible degradation (Attatched photos).
I jiggled the signal cable from the demodulator Q_out to the MC servo. The openloop gain seemed fluctuating (increased) based on the cabling.
I decided to repair these cables by adding solder on the shield.

Even after the repair, the open loop TF didn't show any improvement. I checked the LO level and found that it was -16.7dBm.

I traced the problem down to the frequency generator unit (T1000461). The front panel of the unit indicates the output power for the 29.5MHz output is 13dBm,
while measurement showed it was 6~8dBm (fluctuating). The T1000461 document describes that there is only a wenzel oscillator inside. Does this mean the oscillatorwas degraded??? We need to open the box.

I was not sure what was the LO level. I naively assumed the input is 0dBm. Reducing the attenuation of the dial on the AM Stabilizer unit from 12dB attn to 0dB.
This made the LO level -3.3dBm.


Later at home, I thought this nominal LO level of 0dBm could have been wrong.

The demodulator circuit (D990511) has the amplifier ERA-5 (G=~20dB) at the input. Between the input and the ERA-5 there is a pattern for an attenuator.
Assuming we have no attenuator, the ERA-5 has to spit out 20dBm. That is too much for this chip. I need to pull out the box to see how much is the nominal LO for this box using an active probe.

This decrease/increase of the LO level affects the WFS demod too. According to D980233-B, the input stage has the comparator chip AD96687, which can handle differential voltage of 5.5V.
Therefore the effect is minimal.

Attachment 1: PDRFsignal_cable.JPG  904 kB  | Show | Hide all | Show all
Attachment 2: Qsignal_cable1.JPG  1.154 MB  | Show | Hide all | Show all
Attachment 3: Qsignal_cable2.JPG  1.074 MB  | Show | Hide all | Show all
Attachment 4: Qsignal_cable3.JPG  1.077 MB  | Hide | Hide all | Show all
Qsignal_cable3.JPG
ELOG V3.1.3-